EVA Report – February 24th

Crew 222 EVA Report 24-02-2020
EVA #9
Author: Valentin Prudhomme (Engineer)
Purpose of EVA: Replacement of the battery of MegaAres.
Start time: 4:44pm
End time: 5:13pm

Narrative: Everything went just fine. We rapidly changed the battery and
assure the waterproofness of the experiment, despite the mud all around
the experiment.

Destination: Experiments Site.
Coordinates (use UTM NAD27 CONUS): 12S 0518494-4250679
Participants: Valentin Prudhomme (Engineer, EVA Leader), Bleuenn Roiron
(GreenHab Officer)
Road(s) and routes per MDRS Map: Entrance road – Cow Dung Road
Mode of travel: Rover – Curiosity

Operations Report – February 24th

Crew 222 Operations Report 02/24/2020

SOL: 8

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Friday generator:
Off at 6:30am
On at 9:00pm SOC was 83%

Saturday generator:
Off at 6:00am
On at 9:00 pm SOC was XX%

Diesel Reading – 74%

Station Propane Reading – 29% (refill asked)

Ethanol Free Gasoline – 0.5 gallons

Water (loft tank): 48 gallons

Water Meter: 149453.58 units

Water (static tank) – 304 gallons

Static to Loft Pump used – yes

Water in GreenHab – 72 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no
Hours: 146.2
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover: no
Hours: 132.4
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Opportunity rover used: no
Hours: 69.9
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Curiosity rover used: yes
Hours: 138.5
Beginning charge: 100%
Ending charge: 94%
Currently charging: yes

Notes on rovers: N/A

ATV’s Used: yes
Reason for use: Run during the engineer EVA for the battery.
Oil Added? no
ATV Fuel Used: N/A
Hours the ATVs were Used today: 10 minutes
Notes on ATVs: The second one from the Hab doesn’t have anymore battery.
The red one was very difficult to start but I manage to do it.

HabCar used and why, where? The habcar is now in the shop awaiting repairs.

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: All ok.

Summary of suits and radios: Suits 8 and 9 used and radio tested for the
EVA. all ok.

Summary of Hab operations: N/A

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: After the rain of yesturday, the RAM wasn’t
waterproof and leaked on our electronics.

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: N/A

Sol Summary – February 24th

Crew 222 Sol Summary Report 24-02-2020
Summary Title: The beginning of a new week on Mars
Author’s name: Benjamin Auzou
Mission Status: all things running well
Sol Activity Summary:
This morning was a very hard awakening, with the return of the matinal
workout session. After a long breakfast with pancakes, we worked on
experiments : music for plants, spacesuit, water monitoring. In the
afternoon we started the second part of our human factors experiment,
based on a bomb defusal game. The goal is to study the way we cooperate
to solve a given problem.

Look Ahead Plan:
Tomorrow will be very similar to today. Also we hope to be able to go
outside on an EVA to explore and set the LOAC (one of our experiments)
back on the field.

Anomalies in work: none
Weather: windy
Crew Physical Status: all members safe and sound, the shoulder pain of
our crewmember is gone
EVA: Battery replacement of MegaAres, very windy
Reports to be filed: Journalist report, Photos Report, Photo of the day,
GreenHab Report, Astronomy Report, EVA Report, EVA Request
Support Requested: Trashes in the engineering airlock

GreenHab Report – February 24th

Crew 222 GreenHab Report – 24/02/2020

GreenHab Officer : Bleuenn ROIRON

Environmental control : Both heating and ambiant, door open in the
afternoon( window on the door because of the wind outside) to maintain
temperatures between 20 and 25 °C

Average temperatures: 23°C

Hours of supplemental light: NA

Daily water usage for crops: 8 gallons

Daily water usage for research and/or other purposes: 0 Liters

Water in blue Tank: 72 gallons

Time(s) of watering crops: 2

Changes to crops: NA

Narrative:Today, I started as usual by exposing watercress to UV and
music, and then watered plants. This afternoon I went to harvest some
herbs for our pizzas and some carrots that were really good last time.
Our cucumber grow well and I hope we will eat it in a few days and
Spirulina looks really great in one of the aquarium. I went watering
plants again tonight.

Basil: 18g
Sun carrots: 100g(14 little carrots with leaves)
Chives: 7g
Sage: 1g
Lemon balm: 1g
Oregano: 3g

Support/supplies needed: NA

I am sorry for the mistake of yesterday, everything works very well
about environmental control.

Journalist Report – February 24th

Crew 222 Journalist Report 24Feb2020
Author: Marie Bochard, Crew Journalist
Sol 8
Title: Cold wind blows

The wind today has been very loud. There is no visible movement outside
the windows, because there is no vegetation that bends and moves in the
wind. There is only sand and hills that stay still even when the wind
blows. We were impressed by the sheer speed of it: from our weather
station, we recorded a maximum speed of more than 15 m/s! This is rather
good news because wind dries out the mud that the snow created.
Hopefully, we will have dry ground soon, because going on EVA in mud is
very slippery and dangerous.

Today in the GreenHab, Bleuenn showed us a cucumber that’s almost ready
to be harvested. It’s going to be amazing to have a fresh, crunchy
vegetable! Her spirulina experiment is going well, the algae is growing
and will be harvestable in a few days hopefully. Nina’s aerospring moss
is also growing: out of the four plants she chose, three have grown
significantly and she will be able to transfer the moss into the
vertical garden in a few days. All this is very promising!

For a human factor experiment, we played a game called Keep Talking and
Nobody Explodes, about diffusing a bomb. One person has the bomb on
their computer, and they have to work in collaboration with an expert
who has the manual to diffuse the bomb. We are almost becoming competitive!

We ended the day with a delicious spinach pie staring a stunning “Crew
222” decoration by Nina, and watched Kaamelott all together to have a
good laugh before bed.

Astronomy Report – February 24th

Astronomy Report

Name: Eishi Kim
Crew: 222
Date: 2/24/2020


Robotic Telescope Requested (choose one) MDRS-14

Objects to be Imaged this Evening: Keeping on with the research project,
the following targets are added to the queue for the second round, in
addition to the pending observations: NGC 2683, NGC 3031, NGC 2655, NGC
3344, NGC 4449, NGC 4254, NGC 3338, NGC 2775, M100. This will provide me
with standard images that can be compared to the existing reference
images, in order to try detecting a supernova.

Images submitted with this report: None, but the first standard images
are being analysed using the Photoshop overlay function. No promising
finding is to be reported so far.

Problems Encountered: Frames of NGC 4618, NGC 3521, NGC 4725 and NGC
3675 had WCS registration errors which make them unusable for analysis.
New observations are however planned in the coming days.


Solar Features Observed: N/A

Images submitted with this report: N/A

Problems Encountered: N/A

Copyright © The Mars Society. All rights reserved. | Main Site